OpenStudy (anonymous):

make each fraction equivalent to one-half 7/? = ?/42

OpenStudy (anonymous):

numerically what does one-half look like?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

7/? =7/14 because 7/7=1 and 14/7=2 so it would be 1/2 ?/42 is 42divided by 2 and u get 21 so it is 21/42 because 21 is 1/2 of 42

OpenStudy (anonymous):

so what i would do now is say i have 1/2 right? i know the top part of the first one is 7 right? So what if multiplied the top and bottom by 7?

OpenStudy (anonymous):

what would that look like if i did that?

OpenStudy (anonymous):

1 7/7=1

OpenStudy (anonymous):

no no \[7/?= 1/2 \] you divide both top and bottom by 7 and get 1/2 \[?/42 =1/2\] \[42/2\] (to get 2 parts for the 1/2) \[?/42=21/42 \] because 21 is 1/2 of 42

OpenStudy (anonymous):

OOOOH I get it. You Divided by 2 cause of half

OpenStudy (anonymous):

yes exactly sorry for making it sound so difficult

OpenStudy (anonymous):

so you had the right idea you went 1 x 7 right but to make it so you multiply by 1 and dont change the problem you should multiply by 7/7 so whats 1/2 * 7/7?

OpenStudy (anonymous):

7/14= 1/2*7/7

OpenStudy (anonymous):

good so what about using that method to do the other one?

OpenStudy (anonymous):

21/42 since you dont have the numerator in this problem you MUST divide the denominator by 2 to get 2 halves and 42/2= 21 so 21 is 1/2 of 42